A bullet of mass 3.3 g strikes a ballistic pendulum of mass 1.3kg. The center of mass of the pendulum rises a vertical distance of6.6 cm. Assuming that the bullet remains embedded in the pendulum,calculate the bullet's initial speed.
Answer #1
Momentum initial = momentum final
bullet is mass 1 (m1) and velocity 1(v1)
Pendulum is mass 2 (m2) and Velocity 2(v2)
the bullet sticks in the pendulum after collision there isonly one velocity with the two masses for the final momentum. But since you're not given the final velocity you'll have tocalculate it based on conservation of energy methods ΔKE =ΔPE
The trick is treating this part seperately, set initial PE to0 and initial KE to 0.
Thus KEf = PEf
.5*(m1+m2)*v2 =(m1+m2)*g*d
m1=3.3g = 3.3*10-3kg   but ifyou look closely above masses cancel
d = 6.6 cm = 6.6*10-2 m
So basically you get an equation v= √(2*g*d)
v=√ (2*9.81*.066) = √1.29492 = 1.1379 m/s this isfinal velocity in the next part
so now look at the momentum
m1v1 + m2v2 =(m1+m2)vf      v2= 0 to begin with
m1v1=(m1+m2)vf
.0033v1 = (.0033+1.3)*1.1379  
v1 = 1.4827/.0033
V1 = 449.30 m/s this is initial bullet velocityjust before it hits the pendulum